A merchant gains 15% by selling an article for birr 150;by how much does he sell it to double his profit.

Answer by ikleyn(52925)

.
The condition says that


1.15*x = 150,


where x is the cost of the article for the merchant.


Find x, and then evaluate 1.3*x.


---------------
On percentage problems, see the lessons
    - Percentage problems
    - Percentage word problems (Type 1 problems, Finding the Part)
    - Percentage word problems (Type 2 problems, Finding the Rate)
    - Percentage word problems (Type 3 problems, Finding the Base)
    - More complicated percentage problems
    - Problems on percentage that lead to unexpected results
in this site.

=============
Comment from student: Thank you very much ,but t couldn't understand this "1.15*x = 150" and this "1.3*x". could you please show me by steps?
=============


My response:   1.15 = 1 + 0.15.

                        1.15*x   is the price the merchant sold the article to get 15% profit.

                        1.3*x   is the price to sell in order for to get the profit 30%, which is doubled 15%.


You DEFINITELY need to read the lessons from my list.
Start from the VERY FIRST lesson.